OpenStudy (anonymous):

Determine which equation below has a solution of a = 2. A. 2a + 3a – 6 = –21 B. 8 = 3 + 4a – 3 – 5a C. –2 = 2 + 5a – 2 – 6a D. –4a + 3a – 6 = –4

mathslover (mathslover):

Insert a = 2 in each equation and check , if you get 0 = 0, then that is the answer.

mathslover (mathslover):

Or solve each equation by solving for a. Example : A. 2a + 3a - 6= -21 \(\color{blue}{2a + 3a } = 5a\) Therefore : \(\color{blue}{2a + 3a} - 6 = -21 \\ \implies \color{blue}{5a } - 6= -21\) Add 6 both sides. \(5a \color{red}{-6} + \color{red}{6} = -21 + \color{red}{6}\) Notive : - 6 + 6 there. -6 + 6 = 0 Therefore: \(5a = -21 + 6\) \(5a = -15\) Divide both sides by 5 \(\cfrac{5a}{5} = \cfrac{-15}{5}\) \(a = -3\) But, we have to find that equation which has soln of a = 2 , but here a= -3 . Therefore A.) can not be out correct choice. Similarly do for all.

mathslover (mathslover):

You can also do that by inserting a =2 in each equation... Example : A.) 2a + 3a - 6 = -21 2(2) + 3(2) - 6 = -21 4 + 6 - 6 = -21 4 +0 = -21 4 = -21 By putting a = 2 in the equation, we get 4 = -21 which is not true. Therefore, a = 2 is not the solution of this equation. i.e. A.) is not our correct choice.